Solve for this equation:
7-5(a+1)=8+a

Answer by deborabr(181)

7-5(a+1)=8+a
7-5a-5 = 8+a
-5a-a = 8-7+5
-6a = 6
-a = 6/6
-a = 1 *(-1) to change the signal
a = -1